# BSTcontrol

### v2

@@ 1 Introduction
The advent of sixth-generation (6G) and beyond communication systems indicates a paradigm shift in wireless communications, envisioning a future characterized by unprecedented levels of intelligence, efficiency, and seamless connectivity [1, 2, 3]. To realize this ambitious vision and to navigate the escalating complexity of future wireless networks, novel technological paradigms are urgently needed. Among these, wireless large AI model (WLAM) emerges as a pivotal technology, holding the potential to fundamentally reshape wireless communications [4]. Distinguished by its sophisticated architectures and parameters at massive scales, WLAM offers unparalleled capabilities in data processing, inference, and decision-making, specifically tailored for the unique challenges and opportunities of wireless environments. By leveraging its inherent adaptability and generalization capabilities, WLAM can move beyond current narrow applications, offering the promise of establishing truly AI-native wireless networks capable of handling multifaceted tasks and evolving demands [5].
To elaborate on this vision, AI-native 6G and beyond refers to future communication systems where AI is not merely an add-on feature but is fundamentally integrated into the core design, operation, and optimization of the network fabric. Unlike previous generations where AI might address specific tasks or optimize isolated functions, AI-native networks intrinsically leverage AI across all scenarios and designs. This deep integration aims to enhance wireless performance through AI applications such as optimization and management in physical, network and semantic layers, while utilizing the wireless infrastructure to efficiently support AI operations by enabling edge intelligence and supporting AI security. This holistic and AI-native approach targets unprecedented levels of network automation, self-optimization, resilience, dynamic resource allocation, and the delivery of highly personalized and context-aware services, defining the capabilities expected for the 6G and beyond.
@@ 1.1 Background
Wireless communication has become an indispensable part of modern society, supporting a vast array of applications ranging from personal devices to critical infrastructure. Moreover, large-scale AI is rapidly transforming industries by providing advanced solutions to complex problems. The convergence of these two powerful domains presents a transformative vision for the future, where wireless networks become not only more interconnected but also significantly more intelligent, efficient, and adaptive.
@@
Conventionally, AI models in wireless networks were designed for specific tasks and scenarios, making them highly dependent on extensive data collection and customized training for each application. While effective in isolated cases, these conventional approaches often struggled in dynamic and resource-constrained environments due to their rigidity and lack of scalability [5]. To address these limitations, WLAM have emerged as a groundbreaking solution. WLAM refers to large-scale AI models that seamlessly integrate with wireless communication systems, enhancing network performance through advanced intelligence levels. Unlike conventional models, WLAM can be repurposed for various wireless applications with minimal retraining, leveraging techniques such as prompt engineering and fine-tuning [6]. This flexibility allows WLAM to adapt to evolving user requirements and emerging wireless technologies without undergoing exhaustive retraining cycles. By integrating WLAM, the future of wireless systems moves toward AI-native networks capable of autonomously managing intricate tasks, optimizing performance, and dynamically responding to changing communication landscapes. This advancement paves the way for next-generation intelligent wireless services, offering enhanced efficiency, robustness, and scalability across a wide range of applications.
@@
Large AI models and wireless communications are increasingly synergistic, propelling technological advancement. Large AI models enhance wireless communication systems by processing vast data and learning complex patterns. In return, advancements in wireless communications provide the infrastructure needed to support the computational and data demands of these models, creating a mutually beneficial relationship that fuels innovation. Large AI models improve wireless communications by offering adaptable, generalizable frameworks that outperform traditional task-specific models. They optimize critical functions like network management and signal processing while integrating diverse data sources for better decision-making through multi-modal capabilities [7]. In semantic communication (SemCom), these models boost efficiency by transmitting only essential information, reducing bandwidth needs for future networks [8]. Conversely, wireless communication technologies enable large AI models by addressing their resource demands. Edge intelligence deploys models closer to data sources, cutting latency and energy use through distributed computing [9]. Federated learning supports decentralized training across devices, preserving privacy and reducing communication costs [10]. Over-the-air computation further enhances efficiency by merging communication and computation, minimizing data aggregation overhead [11]. This interplay promises to transform both domains. Large AI models deliver smarter, more efficient wireless solutions, while advanced communication systems empower the development and deployment of these models. Their integration is key to unlocking the potential of future networks and AI applications.
@@
DDIM Denoising Diffusion Implicit Model PPO Proximal Policy Optimization
@@ 1.2.1 Delay
Delay refers to the total time required from the start of local training on the client to the final acquisition of the updated global model. This includes both the local computation time and communication delays. Minimizing delay is essential for improving training efficiency and model performance. A careful balance between computation and communication is necessary to optimize system responsiveness and performance [12].
@@ 1.2.2 Energy Efficiency
Energy efficiency is crucial in WLAM, as energy consumption can impact the sustainability and scalability of the system. This includes both the energy consumed for local computation during model training and for data transmission between clients and the central server. It is vital to design systems that optimize energy usage, balancing between computational load and communication needs. Reducing energy consumption is essential, particularly in resource-constrained environments [13, 14].
@@ 1.2.4 Massive Connectivity
Wireless systems must manage the communication needs of a large number of distributed devices. The system should be designed to handle large-scale connectivity, ensuring that communication between devices remains efficient even as the number of devices increases. Addressing connectivity issues without introducing significant delays is crucial for maintaining system efficiency as the number of devices grows [15, 16].
Table 2: Comparison of our work with existing related surveys
@@
[17]
[18]
[19]
[20]
[5]
[6]
[21]
[22]
@@ 1.2.5 Privacy
Privacy protection is fundamental in WLAM. The system should be designed with privacy-preserving techniques to ensure that sensitive data is not exposed during model training and updates. Ensuring privacy while maintaining performance is an ongoing challenge in WLAM systems [23, 24].
@@ 1.2.6 Generalization
Generalization refers to the ability of a model to perform well on new, unseen data. It is crucial for large AI models to be designed with good generalization ability, ensuring that the model does not overfit to the training data but can adapt to new environments and conditions. Effective designs must focus on creating models that can adapt to various contexts and maintain high performance across different scenarios [25, 26].
@@ 1.3 Motivation and Outline
The structure of this survey, visually depicted in Fig. 1, is organized as follows: Section II lays the foundational groundwork by elucidating the core principles of WLAM, specifically examining the synergistic interplay between large AI models and wireless network architectures. Sections III and IV then delve into the dual facets of WLAM integration, respectively exploring the application of large AI models to enhance wireless communications and the role of wireless communications in enabling large AI models, encompassing architectural considerations and illustrative application scenarios. Section V broadens the scope to investigate the convergence of WLAM with other emerging technologies. Section VI critically analyzes the overarching challenges and outlines promising future research trajectories for WLAM. Finally, Section VII concludes this survey.
@@ 2 Fundamentals of Large AI Models
Large AI models, characterized by their massive scale in parameters, extensive training datasets, and sophisticated architectures, represent a significant leap in AI capabilities. Understanding their core principles is essential for leveraging their power across various domains, including the increasingly complex landscape of wireless communications. These models are not simply larger versions of their predecessors, their scale often leads to qualitatively different behaviors and capabilities [27], necessitating a foundational understanding of their components and development processes.
This section outlines the core components of large AI models and inspirations on WLAM, as summarised in Fig. 3. We begin with basic architectures such as encoder, decoder, and encoder decoder models. Next, we review the lifecycle steps including data collection, pre-training, fine-tuning, and inference. We then examine multimodal alignment to support comprehensive understanding in 6G scenarios. Finally, we discuss deployment strategies that tailor large AI models to the specific requirements of next generation wireless networks.
Figure 3:   The outline of Section II.
@@ 2.1 Model Architecture
The choice of model architecture is crucial in determining how well a system can process, interpret, and generate data. Different architectures are designed to handle specific types of data, such as spatial, temporal, or sequential information, and to optimize performance for tasks like signal processing, resource allocation, and network optimization. In this subsection, we explore various model architectures, ranging from dense networks to specialized approaches like transformers, CNNs, and diffusion models, all of which offer unique advantages in wireless communication systems.
@@ 2.1.1 Dense Network
Dense (fully-connected) neural network is the most basic architecture where each neuron in one layer connects to all neurons in the next [28]. Though simple, dense network provides the foundation for learning arbitrary mappings. However, it is generally less efficient for spatial or sequential data common in wireless signals. Moreover, scaling parameters in dense network can lead to increased computational complexity and longer training times, especially with large datasets common in wireless applications.
@@ 2.1.2 Convolutional Neural Network
Convolutional neural network (CNN) is a class of deep learning models that have proven highly effective for tasks involving spatial data, such as image and signal processing. CNNs use convolutional layers that apply a series of filters to the input data, capturing local patterns and hierarchies of features [29]. This structure makes CNNs suited for processing and analyzing complex patterns in wireless signals, such as those found in channel estimation and signal classification. By leveraging the spatial locality and translational invariance properties of CNNs, wireless communication systems can significantly enhance their performance in tasks like interference mitigation, beamforming, and modulation recognition. Increasing the number of parameters in CNNs, such as depth or number of filters, can improve their ability to learn complex features but may also slow down training convergence and increase the risk of overfitting, especially with limited or biased data from dynamic wireless environments. The computational load during inference also increases with larger CNNs, potentially leading to higher latency under dynamic wireless conditions.
@@ 2.1.3 Transformer
Transformer [30], initially developed for natural language processing, has revolutionized the field of AI with its ability to handle sequential data through self-attention mechanisms. Unlike CNN, transformer does not rely on convolutions or recurrence, allowing them to capture long-range dependencies and relationships in the data. This characteristic is particularly beneficial for wireless communication [31], where understanding the temporal dynamics and dependencies between transmitted and received signals is crucial. Transformers can be employed for tasks such as sequence prediction, channel state information (CSI) feedback, and end-to-end communication system design [32, 33]. While transformers may exhibit poor performance on a small scale due to high computational and memory requirements, they demonstrate outstanding scalability. As the scale of the system increases, their performance improves significantly, making them highly effective for large-scale wireless communication networks.
@@ 2.1.4 Encoder and Decoder Related Architectures
Encoder-only architectures are designed to encode a complete input sequence into a rich, contextualized representation, leveraging attention mechanisms to incorporate bidirectional context. An example is bidirectional encoder representations from transformers (BERT) [34], which masks portions of the input and learns to reconstruct them using the surrounding context. This strategy enables the model to develop deep semantic understanding of the input sequence, making it exceptionally effective for discriminative tasks. In wireless systems, encoder-only models are naturally suited to interpret structured data or signals where global context matters. For instance, such models can be employed to analyze CSI across multiple antennas or subcarriers, identifying spatial and frequency-domain patterns to detect anomalies, perform user classification, or conduct package loss detection [35].
@@
Decoder-only models (e.g., generative pre-trained transformer (GPT) series [36, 37, 38] and LLaMA series [39, 40, 41]) perform autoregressive generation and excel at sequence prediction. They are well-suited for forecasting future wireless traffic patterns or beamforming sequences in time-varying environments. Their in-context learning capabilities also enable real-time adaptation without retraining. The basic principle underlying GPT models is to compress the world knowledge into the decoder-only transformer model by language modeling, such that it can recover (or memorize) the semantics of world knowledge and serve as a general-purpose task solver [42]. Compared to GPT-3, LLaMA incorporates several specific enhancements to maintain similar performance while significantly reducing the number of parameters [39]. For example, in order to enhance training stability, LLaMA normalizes the input of each sub-layer instead of normalizing the output. However, LLaMA cannot generate responsive text [39], and extra fine-tuning is still required.
@@
Encoder-decoder models combine the strengths of both encoders and decoders, offering a two-stage pipeline. The encoder first processes the input into a compressed latent representation, and the decoder subsequently generates the target output conditioned on this representation. Models like text-to-text transfer transformer (T5) [43], bidirectional and auto-regressive transformers (BART) [44], and transformer-based sequence-to-sequence architectures exemplify this class. These models are effective for tasks involving input-output transformations, such as translation, summarization, and question-answering. This structure aligns with tasks like end-to-end communication system modeling or semantic data compression, where input-output transformations are needed. However, they face challenges in scaling and efficiency, especially with longer inputs.

### v4

@@ 1 Introduction
The advent of sixth-generation (6G) and beyond communication systems indicates a paradigm shift in wireless communications, envisioning a future characterized by unprecedented levels of intelligence, efficiency, and seamless connectivity [dang2020should, 10054381, you2021towards]. To realize this ambitious vision and to navigate the escalating complexity of future wireless networks, novel technological paradigms are urgently needed. Among these, wireless large AI model (WLAM) emerges as a pivotal technology, holding the potential to fundamentally reshape wireless communications [shahid2025largescaleaitelecomcharting]. Distinguished by its sophisticated architectures and parameters at massive scales, WLAM offers unparalleled capabilities in data processing, inference, and decision-making, specifically tailored for the unique challenges and opportunities of wireless environments. By leveraging its inherent adaptability and generalization capabilities, WLAM can move beyond current narrow applications, offering the promise of establishing truly AI-native wireless networks capable of handling multifaceted tasks and evolving demands [chen2024big].
To elaborate on this vision, AI-native 6G and beyond refers to future communication systems where AI is not merely an add-on feature but is fundamentally integrated into the core design, operation, and optimization of the network fabric. Unlike previous generations where AI might address specific tasks or optimize isolated functions, AI-native networks intrinsically leverage AI across all scenarios and designs. As envisioned in the comprehensive overview of 6G AI [cui2025overview], this deep integration aims to enhance wireless performance through AI applications such as optimization and management in physical, network and semantic layers, while utilizing the wireless infrastructure to efficiently support AI operations by enabling edge intelligence and supporting AI security. This holistic and AI-native approach targets unprecedented levels of network automation, self-optimization, resilience, dynamic resource allocation, and the delivery of highly personalized and context-aware services, defining the capabilities expected for the 6G and beyond.
@@
Figure 1: The synergistic relationship between Large AI for Wireless and Wireless for Large AI. The top half illustrates the top-down optimization where AI acts as the network brain, while the bottom half depicts the bottom-up infrastructure support provided by the wireless network for AI training and inference.
@@ 1.1 Background
Wireless communication has become an indispensable part of modern society, supporting a vast array of applications ranging from personal devices to critical infrastructure. Moreover, large-scale AI is rapidly transforming industries by providing advanced solutions to complex problems. The convergence of these two powerful domains presents a transformative vision for the future, where wireless networks become not only more interconnected but also significantly more intelligent, efficient, and adaptive [10750803].
@@
Conventionally, AI models in wireless networks are designed for specific tasks and scenarios, making them highly dependent on extensive data collection and customized training for each application. While effective in isolated cases, these conventional approaches face fundamental limitations that hinder their scalability. First, the task-specific paradigm creates isolated intelligence silos, where a model trained for channel estimation cannot generalize to beamforming, leading to excessive management overhead. Second, these models suffer from poor generalization in dynamic wireless environments, requiring frequent and costly retraining when deployment scenarios change. Third, conventional AI lacks the capability to process multi-modal data or interpret high-level human intents, restricting the network from realizing true intent-based autonomy. To address these systemic bottlenecks, Wireless Large AI Models (WLAMs) have emerged as a groundbreaking solution. WLAM refers to large-scale AI models that seamlessly integrate with wireless communication systems, enhancing network performance through advanced intelligence levels. Unlike conventional models, a WLAM serves as a versatile foundation model that can be repurposed for various wireless applications with minimal adaptation [chen2024big]. By leveraging techniques such as prompt engineering and fine-tuning , WLAMs can adapt to evolving user requirements and emerging wireless technologies without undergoing exhaustive retraining cycles [bariah2024large]. This shift paves the way for next-generation intelligent wireless services, offering enhanced efficiency, robustness, and scalability across a wide range of applications.
@@
Large AI models and wireless communications are forming a deeply synergistic ecosystem, propelling technological advancement through a bidirectional interplay. As illustrated in Fig. 1, this relationship can be categorized into two distinct yet interconnected paradigms: “Large AI for Wireless” and “Wireless for Large AI.” In the top-down paradigm of Large AI for Wireless, the AI model functions as the intelligent “brain” of the network. By offering adaptable, generalizable frameworks that outperform traditional task-specific models, WLAMs optimize critical functions ranging from physical layer signal processing to network management [du2024distributed]. For instance, they enable predictive resource allocation and, in the context of semantic communication (SemCom), boost efficiency by transmitting only essential information to reduce bandwidth consumption [jiang2024large]. Conversely, the bottom-up paradigm of Wireless for Large AI depicts the wireless network as the essential infrastructure supporting intelligence. Advanced wireless technologies address the immense computational and data demands of large models. Edge intelligence moves computation closer to data sources to cut latency [zeng2024implementation], federated learning supports decentralized training while preserving privacy [chen2024role], and Over-the-air computation (AirComp) minimizes data aggregation overhead by merging communication and computation [wang2024over]. This closed-loop interaction where the network supports AI operations and AI in turn optimizes the network is key to unlocking the potential of AI-Native 6G applications.
1.2 Key Performance Indicators
@@
DDIM Denoising diffusion tmplicit model PPO Proximal policy optimization
@@ 1.2.1 Delay
Delay refers to the total time required from the start of local training on the client to the final acquisition of the updated global model. This includes both the local computation time and communication delays. Minimizing delay is essential for improving training efficiency and model performance. A careful balance between computation and communication is necessary to optimize system responsiveness and performance [yang2020delay].
@@ 1.2.2 Energy Efficiency
Energy efficiency is crucial in WLAM, as energy consumption can impact the sustainability and scalability of the system. This includes both the energy consumed for local computation during model training and for data transmission between clients and the central server. It is vital to design systems that optimize energy usage, balancing between computational load and communication needs. Reducing energy consumption is essential, particularly in resource-constrained environments [kim2023green, zhao2023energy].
@@ 1.2.4 Massive Connectivity
Wireless systems must manage the communication needs of a large number of distributed devices. The system should be designed to handle large-scale connectivity, ensuring that communication between devices remains efficient even as the number of devices increases. Addressing connectivity issues without introducing significant delays is crucial for maintaining system efficiency as the number of devices grows [yang2020federated, shi2020joint].
Table 2: Comparison of our work with existing related surveys.
@@
[shen2024large]
[lin2023pushing]
[xu2024when]
[tarkoma2023ai]
[chen2024big]
[bariah2024large]
[xu2024large]
[zhou2024large]
@@ 1.2.5 Privacy
Privacy protection is fundamental in WLAM. The system should be designed with privacy-preserving techniques to ensure that sensitive data is not exposed during model training and updates. Ensuring privacy while maintaining performance is an ongoing challenge in WLAM systems [dwork2014algorithmic, kalapaaking2023blockchain].
@@ 1.2.6 Generalization
Generalization refers to the ability of a model to perform well on new, unseen data. It is crucial for large AI models to be designed with good generalization ability, ensuring that the model does not overfit to the training data but can adapt to new environments and conditions. Effective designs must focus on creating models that can adapt to various contexts and maintain high performance across different scenarios [soldati2024design, al2020generalizing, zhurobustnetwork].
@@ 1.3 Motivation and Outline
The structure of this survey, visually depicted in Fig. 2, is organized as follows: Section II lays the foundational groundwork by elucidating the core principles of WLAM. Sections III and IV then delve into the dual facets of WLAM integration, respectively exploring the application of large AI models to enhance wireless communications and the role of wireless communications in enabling large AI models, encompassing architectural considerations and illustrative application scenarios. Section V broadens the scope to investigate the convergence of WLAM with other emerging technologies. Section VI critically analyzes the overarching challenges and outlines promising future research trajectories for WLAM. Finally, Section VII concludes this survey.
@@ 2 Fundamentals of Large AI Models
Large AI models, characterized fundamentally by their massive scale in parameters, extensive training datasets, and sophisticated architectures, represent a significant leap in AI capabilities. In the context of this survey, we specifically define WLAMs based on the intrinsic massive scale of the model parameters and training data, rather than the scale of model deployment across network nodes. This foundational capacity empowers WLAMs to inherently encompass multi-modality by fusing heterogeneous data types including radio frequency signals, text, and images, and achieve cross-task learning across diverse wireless tasks such as channel estimation and network optimization without task-specific retraining. Understanding these core principles is essential for leveraging their power across the increasingly complex landscape of wireless communications. These models are not simply larger versions of their predecessors; their scale often leads to qualitatively different behaviors and capabilities [wei2022emergent], necessitating a foundational understanding of their components and development processes.
This section provides a comprehensive overview of large AI model fundamentals, progressing logically from concept to practice, as summarized in Fig. 4. The analysis begins with the foundational model architectures before proceeding through the essential lifecycle of a large AI model, which covers data collection, supervised pre-training, refinement via reinforcement learning and fine-tuning, and finally, inference strategies. Building upon this lifecycle, the discussion explores crucial practical techniques, including multimodal alignment and deployment. The section culminates by synthesizing these diverse technical elements into a cohesive set of key design principles for WLAM systems.
Figure 4:   The outline of Section 2.
@@ 2.1 Model Architecture
The choice of model architecture is crucial in determining how well a system can process, interpret, and generate data. Different architectures are designed to handle specific types of data, such as spatial, temporal, or sequential information, and to optimize performance for tasks like signal processing, resource allocation, and network optimization. In this subsection, we explore various model architectures, ranging from dense networks to specialized approaches like transformers, convolutional neural networks (CNNs), and diffusion models, all of which offer unique advantages in wireless communication systems.
@@ 2.1.1 Dense Network
Dense (fully-connected) neural network is the most basic architecture where each neuron in one layer connects to all neurons in the next layer. Though simple, dense network provides the foundation for learning arbitrary mappings. However, it is generally less efficient for spatial or sequential data common in wireless signals. Moreover, scaling parameters in dense network can lead to increased computational complexity and longer training times, especially with large datasets common in wireless applications.
@@ 2.1.2 Convolutional Neural Network
CNN is a class of deep learning models that have proven highly effective for tasks involving spatial data, such as image and signal processing. CNNs use convolutional layers that apply a series of filters to the input data, capturing local patterns and hierarchies of features. This structure makes CNNs suited for processing and analyzing complex patterns in wireless signals, such as those found in channel estimation and signal classification. By leveraging the spatial locality and translational invariance properties of CNNs, wireless communication systems can significantly enhance their performance in tasks like interference mitigation, beamforming, and modulation recognition. Increasing the number of parameters in CNNs, such as depth or number of filters, can improve their ability to learn complex features but may also slow down training convergence and increase the risk of overfitting, especially with limited or biased data from dynamic wireless environments. The computational load during inference also increases with larger CNNs, potentially leading to higher latency under dynamic wireless conditions.
@@ 2.1.3 Transformer
Transformer [vaswani2017attention], initially developed for natural language processing, has revolutionized the field of AI with its ability to handle sequential data through self-attention mechanisms. Unlike CNN, transformer does not rely on convolutions or recurrence, allowing them to capture long-range dependencies and relationships in the data. This characteristic is particularly beneficial for wireless communication [wang2022transformer], where understanding the temporal dynamics and dependencies between transmitted and received signals is crucial. Transformers can be employed for tasks such as sequence prediction, channel state information (CSI) feedback, and end-to-end communication system design [wu2023vision, wu2024transformer]. While transformers may exhibit poor performance on a small scale due to high computational and memory requirements, they demonstrate outstanding scalability. As the scale of the system increases, their performance improves significantly, making them highly effective for large-scale wireless communication networks.
@@ 2.1.4 Encoder and Decoder Related Architectures
Encoder-only architectures are designed to encode a complete input sequence into a rich, contextualized representation, leveraging attention mechanisms to incorporate bidirectional context. An example is bidirectional encoder representations from transformers (BERT) [devlin2018bert], which masks portions of the input and learns to reconstruct them using the surrounding context. This strategy enables the model to develop deep semantic understanding of the input sequence, making it exceptionally effective for discriminative tasks. In wireless systems, encoder-only models are naturally suited to interpret structured data or signals where global context matters. For instance, such models can be employed to analyze CSI across multiple antennas or subcarriers, identifying spatial and frequency-domain patterns to detect anomalies, perform user classification, or conduct package loss detection [bertwireless].
@@
Decoder-only models (e.g., generative pre-trained transformer (GPT) series [radford2019language, brown2020language, achiam2023gpt] and LLaMA series [touvron2023llama, touvron2023llama2, dubey2024llama]) perform autoregressive generation and excel at sequence prediction. They are well-suited for forecasting future wireless traffic patterns or beamforming sequences in time-varying environments. Their in-context learning capabilities also enable real-time adaptation without retraining. The basic principle underlying GPT models is to compress the world knowledge into the decoder-only transformer model by language modeling, such that it can recover (or memorize) the semantics of world knowledge and serve as a general-purpose task solver [zhao2023survey]. Compared to GPT-3, LLaMA incorporates several specific enhancements to maintain similar performance while significantly reducing the number of parameters [touvron2023llama]. For example, in order to enhance training stability, LLaMA normalizes the input of each sub-layer instead of normalizing the output. However, LLaMA cannot generate responsive text [touvron2023llama], and extra fine-tuning is still required.
@@
Encoder-decoder models combine the strengths of both encoders and decoders, offering a two-stage pipeline. The encoder first processes the input into a compressed latent representation, and the decoder subsequently generates the target output conditioned on this representation. Models like text-to-text transfer transformer (T5) [2020t5], bidirectional and auto-regressive transformers (BART) [bart], and transformer-based sequence-to-sequence architectures exemplify this class. These models are effective for tasks involving input-output transformations, such as translation, summarization, and question-answering. Therefore, this structure aligns with tasks like end-to-end communication system modeling or semantic data compression, where input-output transformations are needed. However, they face challenges in scaling and efficiency, especially with longer inputs.